OpenStudy (anonymous):

The surface area of two similar solids are 340yd^2 and 1,158yd^2. The volume of the larger solid is 1,712yd^2. What is the volume of the smaller solid?

OpenStudy (dumbcow):

the ratio of the areas is 340/1158 = 170/579 thus the ratio of the side lengths is sqrt(170)/sqrt(579) = .5418 the ratio of volumes is the side length ratio cubed = .5418^3 = .159 Volume of smaller solid = (.159)(1712) = 272.37